OpenStudy (anonymous):

how would i solve: Tyler played 4 basketball games the mean was 10 points the median was 12 points what might each of his scores be?

OpenStudy (anonymous):

help

OpenStudy (m):

8, 8 , 8, 16 (8 + 8 + 8 + 8)/4 = 10 = mean (16 + 8) /2 = 12 = median

OpenStudy (anonymous):

OH MY GOODNESS!!! THANK YOU!

OpenStudy (m):

correction: (8 + 8 + 8 + 16)/4 = 10 = mean

OpenStudy (anonymous):

is the /4 a divide by 4?

OpenStudy (m):

yes

OpenStudy (anonymous):

wanna help me with the rest of my homework?

OpenStudy (m):

if i can i will be on and off though

OpenStudy (anonymous):

YAY thank you! ok so question # 4.

OpenStudy (anonymous):

Desire played 6 basketball games the mean was 18 the range was 8 points what might each of her scores be?

OpenStudy (m):

14 14 14 22 22 22 since mean is add all above and divide by 6 which equals 18 range is difference between largest number and lowers which is 22 - 14 = 8

OpenStudy (anonymous):

okay and question 5: 6 students kept track of how much time they spent reading in a week. mean time= 6 hours mode= 7 hourse what might each of their times be?

OpenStudy (anonymous):

m...when u get done will u help me?

OpenStudy (m):

4 4 7 7 7 7 (4 + 4 + 7 + 7 + 7 + 7)/6 = 6 mode is most # occuring which is 7

OpenStudy (m):

i will be on and off from computer but i will try :)

OpenStudy (anonymous):

okay and the final question is: Francisco bought 5 books the mean price was $23.80 the most expensive book costs $30 what might each of the books cost?

OpenStudy (m):

17.60 23.80 23.80 23.80 30.00 add and divide by 5 = 23.80 = mean

OpenStudy (m):

that all?

OpenStudy (anonymous):

THANK YOU SO MUCH! THAT IS ALL!!

OpenStudy (m):

btw what i did on last one was subtract 23.8 from 30 to get 6.2 then i subtracted 6.2 from 23.8 to keep mean same :)

OpenStudy (m):

good luck!

OpenStudy (anonymous):

this is the one i need help on.